31st August 2006 Hoyle Committee Room 10.00 - 13:00 Present: DWE, MJI, JRL, PSB, MR, STH, EGS, RGM Apologies: WJS Agenda ------ 1. Actions from last meeting 2. Comments on WFAU minutes 3. News from IAU GA 4. Upcoming VDMT meeting 5. CASU help saga 6. Data archives update 7. Optical processing update 8. WFCAM news 9. VISTA news 10. AOB Minutes ------- 1. Actions from last meeting JRL has not implemented the fix to the problem with the UKIRT archive (missing header information problem) - ongoing <<<< STH have had more than one meeting to plan the update to the Calibration PSB Plan document (VIS-SPE-IOA-20000-0002). MJI DWE have had a number of meetings and have a draft version of the Plone EGS site ready for reviewing by the rest of the group. DWE will send out STH a reminder of the address. Please send comments to DWE and EGS. <<<< MR MJI emailed JAC to see if it would be possible to exclude engineering data from the data sent to Cambridge. This is being looking into by JAC. PSB has set up the casuhelp email account. This is aliased to the wfcam account and the .procmail diverts all emails to MJI. DWE has placed the casuhelp email address on the web/Plone pages STH has not yet acquired the tiled supernova remnant image. JMI did not have access to the data, so he will now try Albert Z????? - ongoing <<<< MJI have not drawn up a response to the UDS EDR paper - ongoing <<<< MR STH has not investigated the effects of stacking on the 2MASS/WFCAM non-linearity comparison - ongoing <<<< PSB have had many meetings to decide on what documents are necessary for the STH UK review JRL MJI STH attended a meeting regarding the Z filter (with JPE and PCH). There was PSB an agreement to have a filter that looks very much like the WFCAM one. It would be interesting to know if this has been ordered yet. DWE forwarded the Gaia meeting email to the rest of the group. ALL the ADASS 2006 deadline for abstracts has passed. EGS has submitted an abstract. 2. Comments on WFAU minute PSB described some of the transfer tests that have been carried out and is pleased that they have been done. He added that nothing much is happening on the UKLight front, although the Cambridge end is ready to go. DWE raised the issue of CASU web addresses and the need to inform people of any changes when we swap over to the Plone system. There then followed a discussion regarding the transfer period between the HTML and Plone sites. For the first month or so, it was felt that the 2 systems should be running in parallel. The speed of the Plone server was also discussed and it was decided that a dedicated server should be bought. EGS, PSB and JMI will write a specification for this machine. <<<< It was also decided that this machine would host the ING archive. DWE asked for more details regarding the flagging of persistent objects. MJI said that he, SJW and NCH were considering what needs to be done. MJI mentioned that there had been problems in sending a header updating script to WFAU. The reason was that the ROE spam filter was sending it to a black hole without informing either recipient or sender that this had been done. The solution was to send the Perl script gzipped. 3. News from IAU GA DWE said that the IAU trip had been very successful including some very entertaining planet-definition meetings. Another measure of flux for extended images was mentioned in one of the talks attended (aperture matched colours). MJI said he was aware of this measure. The Commission 8 meeting was more interesting than usual since it was dominated by new telescopes and space projects rather than results from outdated equipment. An extensive document explaining the most recent IAU resolutions was mentioned (USNO Circular 179 by Kaplan). Pat Wallace gave a summary of this which will hopefully by available from the Commission 8 website (which we will be hosting for the next 6 years due to DWE being the new vice-president). 4. Upcoming VDMT meeting MJI said that there was a bit of uncertainty about the next VDMT meeting and that it might be held on the 4th or 5th September (or postponed by at least a week). No agenda has been distributed yet. 5. CASU help saga MJI reported on the latest emails received to give a flavour of the type of questions asked. 12 emails had been received in the period since casuhelp had been set up. Each took about 1 hour each to answer (one email reply is still pending). A number of them covered missing data or data not yet transfered to WFAU ("Where is my data?"). Others asked about what the VDFS will provide. There was then a discussion regarding how visible the casuhelp email address was since none of the above emails went directly to casuhelp. DWE was asked to add the casuhelp email address to the bottom of each Plone page. <<<< STH was asked to contact SJW to place this email address on the UKIDSS website and advertize it. <<<< 6. Data archives update AAT: JRL reported that AAT had swapped over to use PostgreSQL and that the automatic updating of the archive had thus broken. JRL will fix this. <<<< ING: JRL said that EGS had done a good job with the ING archive user interface. This covers 1996+ data and is the default interface. If earlier data is required, you are redirected to the old interface. There are still a few things that need fixing in the headers of some of the files with the 1996+ data such as some FILTER and OBS keywords. JRL will write a script for RWA to run on these files. The new interface is almost ready to go live. WFCAM: MR reported that all tapes up to 31 July had been received (ie. all of 06A). The raw archive was also up-to-date. ESO transfers have about one week of data to go before they have completed 06A. All the data has also been processed apart from one night (27 July), which is still going after 6 days of processing. This is because it is part of of the GPS survey and is a very crowded field. RGM pointed out that the WFCAM summit pipeline sometimes had backlog problems because of these fields. JRL said that this was due to the requirement of processing the whole magnitude range. For the VISTA summit pipeline this will not be a problem since the requirement is for DQC only and the magnitude range to be processed will be limited. UKIRT: JRL said that one fix has been carried out after the Debianizing of the host computer. JRL was asked to email AA that all was OK. <<<< [This doesn't tally with what was said in Section 1. I thought that Jim hadn't finished all the fixes to the UKIRT archive.] 7. Optical processing update MJI said that the revised Software Management Plans for the VST surveys have been sent in on 10 August. He has also been processing the 1/2 Tbyte of UVEX data for Paul Groot. EGS will be rereducing some data (see later) [That's all that was said!] 8. WFCAM news Progress with 06A: As mentioned previously, only one night from 06A is left to process. This night, 27 July, would be a good one to stress test any pipeline. MJI said that he was eyeballing each night and then flagging them as ready to copy afterwards. So far, 20 nights have been checked. MR reported that he was reingesting the data into the DQC database (why?). This will take of order a week to complete. Calibration update: STH said that new colour equations have been derived and that these have been applied to the calculation of the photometric zeropoints of all data processed so far (05A, 05B and 06A). DWE asked STH/MJI to email these colour equations to him. JRL said "Me too". <<<< STH reported that SJW is writing a paper and that from changes in the average colours he is finding an RMS colour variation of 2.5% ie. the photometry is good. MJI suggested matching colour-colour diagrams as a more robust method. ZY zeropoint: STH said that since there were no standards for these passbands, extrapolation from 2MASS values using Y-J vs J-K colour equations were needed. However, this was a non-linear colour equation and A0 stars (for example) were offset by 0.1 mag in Y). Thus not all stars would be on the Vega system for these passbands. [No mention of Z - assume that same applies. I also assume that the reference to the "non-linear" bit is due to the dodginess of extrapolating anything other than a linear equation.] PSF fitting: DWE reported that the tests with real data have gone reasonably successfully. More checking is required with the M17 case although Prague and Dubrovnik have got in the way of this. A new modification has been carried out to the standard pipeline astrometry making it as good as PSF fitting across the entire magnitude range for most circumstances. After some discussion it was felt that there was a need to benchmark the software with respect to other software. STH suggested Phil Lucas as a possible person to contact on this point. It was noted that both Paul Hirst and Sandy Leggett have moved from UKIRT to Gemini. 9. VISTA news VIRCAM status: PSB said that a VIRCAM shipping readiness meeting was being held (CASU not invited). The chip that was broken has now been replaced. There was a possibility that we would get a real QE response curve for this chip. He again stressed that if there was another opportunity to acquire some data, it would be very helpful. Apparently there have been some mirror problems which might cause a three month delay. An idea has been floated of installing a 1m mirror in the meantime in order to test the telescope. VDFS docs status: The ETC has now been updated to have the new overhead algorithms. This results in a 10% better efficiency in the observing. JRL said that he wanted to update the figures in the DRLD. STH will provide these. <<<< It was stated that CASU would probably get the 6 month extension to the VISTA development grants. This would be for 3 FTEs and would cover JRL, PSB, STH (0.5) and EGS (0.5). The later would buy out this AstroGrid component. It was also announced that PSB would be the VISTA pipeline manager. The VISTA data transfer to ESO will be using disc drives. These would be standardized to 250 Gbyte Linux file system discs. It was felt that larger discs would be more useful. PSB thought that an ICD might be needed on this issue. He will probably have a chat with the appropriate people at ADASS. 10. AOB ADASS 2007: JRL said that he wanted a meeting next Monday to sort out ADASS 2007. Hardware update: MJI reported that the equipment grant has at last got through so we can pay back the University the money that we've already spent. /data/cass55c has been scrambled and all data lost. This RAID array has now been moved to the APM machines and called /data/apm23c (7 Tbyte RAID5). It is unclear what happened. A new 10 Tbyte RAID6 array has been acquired. This is called /data/apm23d. Power Cut: To prevent any problems from the power cut scheduled for the weekend 16/17 September, MJI & JMI will switch off all APM machines on the Friday. Bringing them all back up may be an interesting experience. Upcoming UK review: MJI said that there will be a need for a pile of documents. The UK Impact Document will be written by JPE. The already written ESO Calibration Plan and DRLD will also be used since the ESO pipeline will be used on the VISTA data. Additionally there will be a UK appendix/extra document to explain tiling. There was some discussion if there was a need for an overview document. Perhaps an updated version of the SPIE paper. MJI will be responsible for this. There will be a need to describe the hardware system. PSB and STH will do this. A need was also felt for a precis of the UK URD, probably drawn up by STH and MJI. A separate Photometric Calibration plan was also going to be written by STH. The formal list of documents needed for the UK VISTA Review (and their authors) is as follows: .......... . . .......... Continuing Actions ------------------ JRL finish the fix to problem with UKIRT archive (missing header information) STH acquire tiled supernova remnant image MJI draw up a response to UDS EDR paper MR STH investigate effects of stacking on 2MASS/WFCAM non-linearity comparison New Actions ----------- DWE remind ALL about location of Plone site etc. ALL review Plone site and send comments to DWE and EGS EGS write a specification for the Plone server PSB JMI DWE add the casuhelp email address to the bottom of each Plone page STH contact SJW to place this email address on the UKIDSS website and advertize it JRL repair AAT archive (PostgreSQL problems) JRL email AA that all is OK with UKIRT archive MJI email new WFCAM colour equations to DWE and JRL STH STH provide JRL with new figures for DRLD